OpenStudy (anonymous):

Let cos p=.22 What is cos p/2?

OpenStudy (harsimran_hs4):

cos 2p = 2cos^2 p - 1 can you proceed now ?

OpenStudy (anonymous):

No

terenzreignz (terenzreignz):

Use this formula, then: \[\Large \cos\left(\frac{p}2\right)=\sqrt{\frac{1+\cos(p)}{2}}\]

OpenStudy (anonymous):

.78 is that right

terenzreignz (terenzreignz):

Approximately, yes.

OpenStudy (anonymous):

Thank you can i put up another one not sure but I think my answer is wrong on another question

terenzreignz (terenzreignz):

Okay, go ahead.

OpenStudy (anonymous):

sin m=.11 what is sin m/2 I got 0.055 but seems to easy that way?

terenzreignz (terenzreignz):

First, you need to get cos m

OpenStudy (anonymous):

I don't know how to ?

terenzreignz (terenzreignz):

Well, you can always use this formula... \[\Large \cos^2 m + \sin^2 m =1\]

OpenStudy (anonymous):

Is .78 the answer for this to?

terenzreignz (terenzreignz):

Let's find out. First, find cos m using that formula I posted...

OpenStudy (anonymous):

I got .78 ? for cos 2 not right is it

OpenStudy (anonymous):

mean cos m

terenzreignz (terenzreignz):

What formula have you been taught for half-angle of sine?

OpenStudy (anonymous):

that my problem not been given one been looking all online for help feel like we a skipping sections in class

terenzreignz (terenzreignz):

Okay, when you are given the sine, to get the cosine, use this formula: \[\Large \cos(m) = \sqrt{1-\sin^2m}\]

OpenStudy (anonymous):

.883176087 ?

terenzreignz (terenzreignz):

nope

OpenStudy (anonymous):

.223606798

terenzreignz (terenzreignz):

Please don't guess :) What is sin m?

OpenStudy (anonymous):

.11 then i divided by 2

terenzreignz (terenzreignz):

No, just sin m for now, not cos m

OpenStudy (anonymous):

0.055 and put in equation you gave

OpenStudy (anonymous):

not sure

terenzreignz (terenzreignz):

sin m is 0.11, yes?

OpenStudy (anonymous):

0.88

terenzreignz (terenzreignz):

No, sin m... it's 0.11

terenzreignz (terenzreignz):

If you have the sine, to get the cosine, use this formula (altered so you can better understand) \[\Large \cos(m) = \sqrt{1-[\sin(m)]^2}\]

OpenStudy (anonymous):

answer is .78